Centripetal force acting on a body rotating in a circle if the body moves in a circle of radius half of the initial value keeping other quantities constant then percentage change in the centipetal force is
A) 300 % B)100% C)200% D)400%
1
Expert's answer
2018-09-03T15:40:08-0400
The initial centripetal force is
F=(mv^2)/r.
The final centripetal force is
F^'=(mv^2)/(r/2)=2F.
The percentage change in the centripetal force is
(F^'-F)/F 100%=(2F-F)/F 100%=100%